Wildcat Ltd, a manufacturing company sold a machinery for Rs 8 lacs at the year end. The company had purchased the machinery four years back for Rs 15 lacs and had depreciated the same using written down value method of depreciation @ 20%. As an accounts executive of Wildcat Ltd, calculate the WDV of the asset for the four years, accumulated depreciation for four years and profit/loss on sale, if any.
1
Expert's answer
2019-11-11T15:44:11-0500

Purchase price of the machinery at Year 1 = 1500000

At the end of Year 1 : -

Depreciation @ 20% on 1500000 = 300000

WDV = 1500000 - 300000 = 1200000

Accumulated Depreciation = 300000


At the end of Year 2 : -

Depreciation @ 20% on 1200000 = 240000

WDV = 1200000 - 240000 = 960000

Accumulated Depreciation = 300000+240000 = 540000


At the end of Year 3 : -

Depreciation @ 20% on 960000 = 192000

WDV = 960000 - 192000 = 768000

Accumulated Depreciation = 540000+192000 = 732000


At the end of Year 4 : -

Depreciation @ 20% on 768000 = 153600

WDV = 768000 - 153600 = 614400

Accumulated Depreciation = 732000+153600 = 885600


Sale price of the machinery = 800000

WDV at the end of year 4 = 614400

Profit on sale of machinery = 185600